OpenStudy (anonymous):

Directions say: Determine whether the lines are intersecting, parallel or the same (check slopes) How many points in common, Give solution if any: 3x-y=2 6x-2y=3 Explain.

OpenStudy (anonymous):

equation of a line: y=mx + b equating your givens: y=3x-2 2y=6x-3 (divide both sides by 2 to eliminate constant 2. so we will get:) y=3x-3/2 lines are parallel.. you can graph it by substituting x. let's say from -3 to 3.. slopes: both slopes are 3.

OpenStudy (anonymous):

I still don't understand.. But thanks!

OpenStudy (anonymous):

what part?

OpenStudy (anonymous):

All of it.. could you draw it?

OpenStudy (anonymous):

ok, wait. this is not easy drawing. xD

OpenStudy (anonymous):

Well it's okay

OpenStudy (anonymous):

did you know why i came up with y=3x-2 from 3x-y=2?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

but then wouldn't be -y=-3x+2 & -2y=-6x+3?

OpenStudy (anonymous):

because if you transfer a number from other side, the sign will change. let's say -2+3=1 if you transfer -2 to the right side, it will be 3=1+2..

OpenStudy (anonymous):

-2+3=1.. if you transfer -2 to the right side, it will be 3=1+2

OpenStudy (anonymous):

the red line is your y=3x-2 and the green line is your y=3x-3/2 and for the slope, from the equation y=mx+b, m is the slope. so the slope for the two linear equations is 3. the rest are in the images. :D http://www.fileden.com/files/2007/10/19/1523433/IMG_0022.PNG http://www.fileden.com/files/2007/10/19/1523433/SAM_1559.JPG

OpenStudy (anonymous):

Thank you sooooo much! you did so much work! Lol Pretty handwriting too!(:

OpenStudy (anonymous):

lol. no problemo.. just trying to help. :D